ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

16. Applications

1991 MS Classification: 94 Karlof, John, University of North Carolina at Wilmington, U. S. A. A Coding Scheme for the Gaussian Multiple Access Channel We present a coding-decoding scheme for the Gaussian multiple access channel (GMAC). This scheme is a generalization of a technique originally developed by Urbanke. Urbanke begins by choosing a linear code over the finite field F_{3} and a set of linearly independent vectors, one for each user. Each user selects an information bit from F_{3} to be transmitted over the GMAC. Each user’s information bit and vector is encoded using the linear code and then mapped to a vector in R^{n}. The vector representing the sum of the mapped codewords plus Gaussian noise is sent to the decoder. We extend Urbanke’s results by allowing the choice of any finite field F_{k}, with k an odd prime.